The sum of two numbers is 101. If four times the smaller number is subtracted from the larger number, the result is 6. Find the two numbers
Answer by ewatrrr(24785)

x the smaller number. Larger 101-x
(101 - x) - 4x = 6
-5x = -95
x = 19, the smaller number. Larger is 82 101-19
And...checking
82 - 76 = 6